Approaching a street with heavy traffic, a jeep slows down and reduces its velocity from 90
km/h to 10 km/h in 10 seconds. Calculate the acceleration of the jeepney.

Answer by amarjeeth123(574)

Initial velocity v=90km/hr=90(1000m/3600sec)=90(5/18)=25m/sec
Final velocity u=10km/hr=10(5/18)=(50/18)=2.7778 m/sec
Time t=10 sec
Acceleration=change in velocity/time=(25-2.7778)/10=22.2222/10=2.22 m/sec^2
Acceleration of the jeepney=2.22 m/sec^2